Determine whether the following series converge or diverge. Fully justify your answer.T(-1)(n? – 2n) 400n3 + 78972 2
Answer #1

Eton (nan) = {an (let) (400 mt 7899) an= EM 3 1 1 - 1 4 = fish ( 12 ) (400+ 780) Thoot 789) nem Jant = nima t o - 1400-789) 4